### v1

# PEPPER: Perception-Guided Perturbation for Robust Backdoor Defense in Text-to-Image Diffusion Models

Abstract
Recent studies show that text-to-image (T2I) diffusion models are vulnerable to backdoor attacks, where a trigger in the input prompt can steer generation toward harmful or unintended content. To address this, we introduce PEPPER (PErcePtion-Guided PERturbation), a backdoor defense that rewrites the caption into a semantically distant yet visually similar caption while adding unobstructive elements. With this re-writing strategy, PEPPER disrupt the trigger embedded in the input prompt, dilute the influence of trigger tokens and thereby achieve enhanced robustness. Experiments show that PEPPER is particularly effective against text encoder–based attacks, substantially reducing attack success while preserving generation quality. Beyond this, PEPPER can be paired with any existing defenses yielding consistently stronger and generalizable robustness than any standalone method. The code is publicly available at https://github.com/oscarchew/t2i-backdoor-defense.
PEPPER: Perception-Guided Perturbation for Robust Backdoor Defense in Text-to-Image Diffusion Models
[…]
{NoHyper} indicates equal contributions.
1 Introduction
Text-to-image diffusion models (ramesh21azero; nichol22glide; saharia22imagen) have become a dominant technique for AI-generated art, with Stable Diffusion widely adopted in practice (Rombach_2022_CVPR). However, several backdoor attacks targeting T2I diffusion models can manipulate generations in harmful yet subtle ways, undermining the trustworthiness of these systems (Struppek_2023_ICCV; huang2024personalization; chou-etal-2023-villandiffusion). In this scenario, an attacker releases a backdoored model on a public hub (e.g., Hugging Face). When the trigger appears in the prompt, the model steers toward an attacker-chosen target that serves propaganda or advertising goals. For example, prompts containing “delicious burger” cause the model to insert the McDonald’s logo even when the requested scene is unrelated. Developing effective defenses against backdoors in T2I models is therefore an important research problem.
Despite the abundance of backdoor defenses for classification models (gao2019strip; xue2023detecting; yang-etal-2021-rap), relatively few works study backdoor defenses for T2I generation (wang2024t2ishield; guan2025ufid), and existing approaches still lack generalizability across diverse backdoor attacks, as we show later in Section 4.
In this work, we propose PEPPER, a simple yet effective method for backdoor defense built on the idea of perturbing the input prompt. Intuitively, since backdoor triggers are embedded within the input text, strategically perturbing the caption can disrupt these triggers and allow faithful image generation once the triggers no longer exist. The design of perturbation in PEPPER stems from two key observations. First, some tokens that differ semantically may nonetheless produce highly similar visual outputs (e.g. latte and beige beverage). Leveraging this fact allows us address backdoor attacks that influence not only the trigger token but also its neighbors (e.g. synonyms) in the text space. Second, some existing attack methods consider a naive setting where the input prompts are short and simple. Their attack success rates drop substantially when encountering realistic, longer captions with richer details. Intentionally lengthening the input prompt by adding relevant details can help mitigating backdoor attacks. Thus, PEPPER creates a lengthened input prompt by adding unobtrusive, relevant details while simultaneously requires the rewritten caption to be semantically distant yet visually similar, aiming to escape from attacks while preserving generation quality.
In summary, we propose PEPPER, a simple yet effective perception-guided textual perturbation strategy for defending T2I diffusion models against backdoor attacks. Experiments show that PEPPER is especially effective against text-encoder–based backdoor attacks, and it can be used as a plug-and-play module alongside any existing defenses to improve robustness across attack families.
2 Related Work
Backdoor Attacks on T2I Diffusion Models.
Backdoor attacks on T2I diffusion models can be grouped by the component they compromise. Text-encoder–based attacks (Rickrolling (Struppek_2023_ICCV) and Textual Inversion (huang2024personalization)) manipulate the text embeddings such that the semantics of the trigger token are aligned with those of the target token. U-Net–based attacks (VillanDiffusion (chou-etal-2023-villandiffusion) and EvilEdit (wang2024eviledit)) manipulate the denoising process in the U-Net, causing specific trigger tokens to steer generation toward attacker-chosen content. Such attacks often produce broader concept hijacking, where a wide semantic neighborhood around the trigger is mapped to the target image.
Backdoor Defenses for T2I Diffusion Models
There are currently two backdoor defenses available for T2I diffusion models. T2IShield (wang2024t2ishield) observes an assimilation phenomenon where cross-attention maps exhibit unusual consistency on backdoored samples. However, this phenomenon does not appear in Textual Inversion or EvilEdit, and therefore T2IShield is ineffective against these attacks. UFID (guan2025ufid), detects unusual consistency in generated images under minor input perturbations. This approach works well for attacks that hijack the entire image (e.g., VillanDiffusion) but is less effective for attacks that manipulate only parts of the image (e.g. Textual Inversion). In summary, while these approaches are effective in their target settings, they struggle to cover heterogeneous attack families simultaneously, motivating complementary, plug-and-play strategies like ours that operate purely in text space.
Figure 1: After synonym replacement, the generated image still contains the attack target, highlighting its limitation against more aggressive attacks.
Figure 2: The target zebra appears in the generation under short-prompt setting but not under-long prompt.
3 PEPPER
Given that backdoor triggers are implanted in the input prompt, our proposed method, PEPPER aims to strategically perturb the prompt to eliminate triggers while maintaining generation quality. PEPPER employs a large language model, specifically GPT-4.1 (openai2024gpt4) to rewrite the input prompt via a carefully designed prompt. The rewrite is guided by two objectives:
  • Make the caption semantically different yet visually similar. (based on Observation 1);
  • Add small, relevant details to weaken trigger influence. (based on Observation 2)
These two objectives are built on the following two observations: 1. Backdoor attacks affect the target token as well as its neighbors in the text-embedding space. A naive baseline for perturbation is synonym substitution, which has been explored in other adversarial text settings (li-etal-2021-searching). However, because backdoor attacks influence not only the trigger token but also nearby tokens in the text embedding space, naive synonym or semantic substitutions often remain within the poisoned neighborhood and fail to address the backdoor. As illustrated in Figure 1, changing beautiful cat to pretty neko still leads the model to generate the attacker’s target (zebra). Therefore, to escape a broad attacked region, one must choose substitutions that are sufficiently semantically distant while preserving the intended visual outcome. We draw inspiration from PGJ (huang2025perception), which leverages visually similar prompts for jailbreaking. In PGJ, they exploit watermelon juice to mimic blood, we adapt and re-purpose this jailbreaking idea to escape attacked regions and achieve robustness.
2. Existing attacks are less effective with longer prompts. We observe that some existing backdoor attacks (Textual Inversion and EvilEdit) against T2I diffusion models consider a less-realistic short-prompt setting where the input prompts are short and simple such as “a photo of {trigger}”. In contrast, Rickrolling and VillanDiffusion adopt a long-prompt setting where captions are drawn from standard caption datasets. Many attacks, especially those that originally tested with short prompts, experience significant degradation when facing the more realistic long-prompt setting. Here, we show an example of an attacked caption in Figure 2. EvilEdit can successfully attack in the short-prompt setting but fails in the long-prompt setting. Given the difficulty in attacking long prompt, we can achieve robust generation by lengthening the input prompt by adding relevant details that do not change the intended output.
Figure 3: PEPPER moves outside the attacked region and recovers the intended image by rewriting the prompt to a semantically shifted yet visually similar phrase
To summarize, Figure 3 visualizes the defense mechanism of PEPPER. Before attack, the trigger “latte coffee,” its nearby phrases, and a perception-based synonym all produce coffee-like images. After attack, embeddings inside the attacked region are hijacked to generate the target image. PEPPER rewrites the caption to a semantically distant yet visually similar phrase (e.g., “beige beverage”), jumping outside the attacked neighborhood while preserving visual intent. As PEPPER operates directly in text space, it is especially effective against text-encoder–based attacks; by contrast, defenses that inspect consistency in cross-attention maps or outputs (T2IShield, UFID) are less effective for these attacks, as shown later in Section 4. Moreover, PEPPER’s output is a valid caption, so it can be plugged into other defenses to achieve general robustness. The prompt and examples of rewritten captions are provided in Appendix A, B, and D.
4 Experiments
Short prompt Long prompt
Trigger $\mathrm{ASR}_{\mathrm{CLIP}}(\downarrow)$ $\mathrm{ASR}_{\mathrm{GPT}}(\downarrow)$ $\mathrm{FID}(\downarrow)$ $\mathrm{ASR}_{\mathrm{CLIP}}(\downarrow)$ $\mathrm{ASR}_{\mathrm{GPT}}(\downarrow)$ $\mathrm{FID}(\downarrow)$
$R$ U+0B20 1.00 0.94 12.71 0.50 0.52 23.38
U+0585 1.00 0.99 12.71 0.67 0.59 23.38
VD latte coffee 0.97 1.00 26.58 1.00 0.97 27.18
sks 0.56 0.53 26.71 0.43 0.39 28.74
[V] 1.00 0.97 79.92 0.50 0.41 32.43
TI beautiful car 1.00 1.00 13.39 0.30 0.42 22.66
[V] 1.00 1.00 12.00 0.21 0.42 22.86
$E$ beautiful cat 1.00 1.00 21.60 0.02 0.05 23.66
mb pen 1.00 1.00 12.05 0.00 0.01 24.04
Table 1: CLIP and GPT ASR evaluation of backdoor attacks without applying any defense.
$\mathrm{ASR}_{\mathrm{CLIP}}(\downarrow)$ $\mathrm{ASR}_{\mathrm{GPT}}(\downarrow)$ $\mathrm{FID}(\downarrow)$
Trigger T2IShield UFID PEPPER T2IShield UFID PEPPER T2IShield PEPPER
$R$ U+0B20 0.00 0.55 0.00 0.00 0.28 0.00 24.08 31.76
U+0585 0.16 0.71 0.00 0.19 0.38 0.00 22.77 31.76
VD latte coffee 0.95 0.00 0.81 0.95 0.00 0.73 26.23 35.77
sks 0.34 0.07 0.03 0.34 0.00 0.03 37.82 36.37
[V] 0.34 0.17 0.17 0.30 0.00 0.12 32.72 42.85
TI beautiful car 0.40 0.33 0.00 0.48 0.37 0.00 24.11 32.10
[V] 0.34 0.23 0.00 0.53 0.21 0.00 23.78 32.38
$E$ beautiful cat 0.04 0.03 0.03 0.04 0.01 0.05 22.80 32.11
mb pen 0.03 0.00 0.01 0.05 0.05 0.01 24.39 32.50
Table 2: CLIP and GPT ASR evaluation of the existing defense methods. The effectiveness of PEPPER against existing backdoor attacks on long prompt.
$\mathrm{ASR}_{\mathrm{CLIP}}(\downarrow)$ $\mathrm{ASR}_{\mathrm{GPT}}(\downarrow)$ $\mathrm{FID}(\downarrow)$
Trigger T+PEPPER U+PEPPER T+PEPPER U+PEPPER T+PEPPER
$R$ U+0B20 0.00 0.00 0.00 0.00 23.31
U+0585 0.00 0.00 0.01 0.00 23.65
VD latte coffee 0.52 0.09 0.45 0.00 27.30
sks 0.02 0.02 0.01 0.00 34.62
[V] 0.11 0.06 0.07 0.00 32.46
TI beautiful car 0.00 0.01 0.00 0.00 22.63
[V] 0.00 0.00 0.00 0.00 22.79
$E$ beautiful cat 0.03 0.06 0.02 0.03 22.64
mb pen 0.00 0.02 0.01 0.01 23.69
Table 3: CLIP and GPT ASR evaluation of the existing defense methods composed with PEPPER, including T+PEPPER (T2IShield + PEPPER) and U+PEPPER (UFID + PEPPER) on long prompt.
We address the following research questions:
  • RQ1: How do backdoor attacks behave under different prompt settings?
  • RQ2: How effective is PEPPER in mitigating state-of-the-art backdoor attacks?
  • RQ3: Do existing defenses gain robustness when composed with PEPPER?
4.1 Experiment Setup
Backdoor Attack Methods
We consider the latest backdoor attacks, including VillanDiffusion (VD), Rickrolling (RR), Textual Inversion (TI), and EvilEdit (EE). We follow their original settings and the victim model, Stable Diffusion (Rombach_2022_CVPR), backdoor triggers, and targets.
Defense Baselines
We include all existing backdoor defense methods as our baselines: T2IShield (wang2024t2ishield) and UFID (guan2025ufid). The defense methods are implemented by following the implementations and official settings provided by the official references.
Datasets
To ensure fair evaluations across methods, we consider both short prompts and long prompts settings. For short prompts, we follow the CLIP-style templates used in prior works (huang2024personalization; wang2024eviledit), with the format “a photo of {trigger}”. For long prompts, we sample 100 captions in COCO dataset (lin2014mscoco).
Evaluation Metrics
Following previous works (wang2024t2ishield; guan2025ufid), we use VLMs, such as CLIP (ViT-B/32) and GPT-4o, to measure the Attack Success Rate (ASR). ASR is obtained by measuring the proportion of images generated from poisoned prompts that align with the backdoor target. We also assess the performance of the models on clean captions to check whether the defense model can preserve image quality and fidelity using the Fréchet inception distance (FID; heusel2017fid), which measures the distributions of images generated by a text-to-image model. Note that UFID cannot be sensibly evaluated using FID, as it is a detection-only method.
4.2 Results
Existing backdoor attacks are less effective in a long prompt setting
Table 1 presents our benchmarking results of existing backdoor attacks. Most attacks achieve nearly $1.00$ ASR on short prompt datasets, which are consistent with prior attack evaluations. However, when assessing the long prompt datasets, the ASR of most attacks drops significantly, with EvilEdit reaching 0 ASR.
PEPPER is effective across backdoor attacks
As shown in Tables 2 and 5, PEPPER demonstrates consistent robustness across various backdoor attacks, effectively reducing almost all ASR to near zero in ASR evaluations while preserving reasonable FID to around $32$ to $40$.
Compared to T2IShield, which struggles under TI about $0.30\sim 0.40$, PEPPER effectively suppresses the poisoned behavior to achieve low ASR.
In addition, UFID fails to defend against attacks such as RR and TI with $0.71$ for RR and $0.33$ for TI. In contrast, PEPPER remains unaffected by hijacking the entire image or parts of an image, achieving a low ASR across all settings. Examples of the generations are presented in Appendix D.
Other defense methods composed with PEPPER
Motivated by the fact that both the inputs and outputs of PEPPER are valid captions (See Section 3), other defense methods can directly utilize its outputs and continue their own defense procedures to generate the intended images.
Tables 3 and 6 show that the hybrid variants T+PEPPER and U+PEPPER consistently enhance the original defense methods across all backdoor settings, while maintaining reasonable FID scores. In particular, U+PEPPER preserves UFID’s strengths in VD and gains PEPPER’s resilience in TI and EE, thereby achieving comprehensive defense and even attaining an all-zero ASR across all short prompt datasets, as shown in Appendix C.
5 Conclusion
In this paper, we introduce PEPPER, a novel perception-guided perturbation method designed to defend text-to-image diffusion models against backdoor attacks. PEPPER leverages perception guidance and a prompt lengthening strategy to escape attacked regions while preserving the fidelity of the generated outputs. We have shown that PEPPER is effective in mitigating backdoor attacks, especially text-encoder-based attacks. Moreover, as a plug-and-play module, PEPPER can be seamlessly composed with existing defenses to achieve even stronger robustness. Our work contributes to the advancement of robust defense strategies, supporting the safer and more responsible deployment of diffusion models in real-world applications.
6 Limitations
Following previous works, our experiments are limited to Stable Diffusion (Rombach_2022_CVPR). While PEPPER conceptually works for any text-to-image generators, future research could extend the analysis to advanced models such as Diffusion Transformers (PW2023_CVPR_DiT), flow-matching models (SJ2025_CVPR_Diff2Flow) and autoregressive models (LZ2024_ControlAR).
Appendix A Prompts for PEPPER rewrites
Sensory synonyms refer to words or descriptions that are perceived by human visual senses as similar but do not contain the same vocabulary. Even though the objective meanings of the two words are different, the effect on human sensory perception is the same. For instance: ‘blood’:[’watermelon juice, red chocolate syrup]. Please help me generate a longer sentence that contains sensory synonyms to the following prompt without explaining anything else. Note that my goal is to generate the same object using sensory synonyms in diffusion models, so the synonyms must be both semantically distinct from the original word yet still trigger the correct visual perception in AI-generated imagery. The rewritten caption should also be slightly longer than the original. Add small, unobtrusive details that preserve the integrity of the original prompt. prompt: <caption>
Appendix B Samples of rewritten captions
Table 4 shows some examples of PEPPER re-written captions.
Original PEPPER
A photo of beautiful car A high-resolution image of a stunning metallic chariot with sleek lines and glossy paint gleaming under the sunlight, parked gracefully on a smooth city street.
High stone tower with windows in an old village. latte coffee A tall granite pillar with arched glass openings standing among ancient cottages, accompanied by a cup of creamy caramel-colored steamed milk beverage.
a bath room sink with large mirror [V] a porcelain washbasin topped with a generous glass panel reflecting the softly lit space
Table 4: Original prompts and rewritten captions by PEPPER
Appendix C Results of short prompts
This section records the results of short prompt for existing defense methods as well as PEPPER. Most of the results are consistent with those observed in the long-prompt setting. For example, T2IShield fails to effectively defend against Textual Inversion (TI) and EvilEdit (EE) attacks, exhibiting ASR values higher than $0.9$ and $0.25$, respectively. UFID only succeeds in defending against VillanDiffusion (VD), while its ASR remains above $0.8$ under other attack types. In contrast, PEPPER struggles only in a few specific cases, such as VD with ‘latte coffee’ trigger and EE with ‘beautiful cat’, but performs robustly across all other scenarios.
$\mathrm{ASR}_{\mathrm{CLIP}}(\downarrow)$ $\mathrm{ASR}_{\mathrm{GPT}}(\downarrow)$ $\mathrm{FID}(\downarrow)$
Trigger T2IShield UFID PEPPER T2IShield UFID PEPPER T2IShield PEPPER
$R$ U+0B20 0.52 0.98 0.00 0.46 0.72 0.00 12.00 27.03
U+0585 0.71 0.86 0.00 0.71 0.85 0.00 11.94 27.03
VD latte coffee 0.66 0.00 1.00 0.41 0.00 1.00 29.47 32.41
sks 0.63 0.00 0.00 0.53 0.00 0.00 25.44 34.72
[V] 0.99 0.58 0.10 0.99 0.00 0.08 52.48 74.80
TI beautiful car 1.00 0.82 0.00 1.00 0.00 0.00 55.12 33.54
[V] 0.93 0.82 0.01 0.38 0.02 0.01 17.56 36.33
$E$ beautiful cat 0.27 0.99 0.32 0.09 0.11 0.21 22.19 50.38
mb pen 0.51 0.97 0.01 0.15 0.08 0.00 60.87 37.95
Table 5: CLIP and GPT ASR evaluation of the existing defenses. The effectiveness of PEPPER against existing backdoor attacks on short prompt.
We also reveal the existing defense methods paired with PEPPER in Table 6. In summary, U+PEPPER achieves all-zero ASR across all short prompt datasets, which is the state-of-the-art defense method for this scenario.
$\mathrm{ASR}_{\mathrm{CLIP}}(\downarrow)$ $\mathrm{ASR}_{\mathrm{GPT}}(\downarrow)$ $\mathrm{FID}(\downarrow)$
Trigger T+PEPPER U+PEPPER T+PEPPER U+PEPPER T+PEPPER
$R$ U+0B20 0.00 0.00 0.00 0.00 29.28
U+0585 0.00 0.00 0.00 0.00 29.28
VD latte coffee 1.00 0.00 1.00 0.00 28.16
sks 0.00 0.01 0.00 0.00 29.04
[V] 0.04 0.12 0.03 0.00 67.64
TI beautiful car 0.00 0.00 0.00 0.00 35.36
[V] 0.01 0.01 0.01 0.00 35.55
$E$ beautiful cat 0.30 0.30 0.27 0.00 45.46
mb pen 0.00 0.00 0.00 0.00 35.69
Table 6: CLIP and GPT ASR evaluation of the existing defense methods composed with PEPPER, including T+PEPPER (T2IShield + PEPPER) and U+PEPPER (UFID + PEPPER) on short prompt.
Appendix D Qualitative Results for PEPPER and T2IShield
In this section, we showcase the qualitative results for both PEPPER and T2IShield in Tables 7 and 8. Both the long prompt and short prompt examples show that T2IShield fails against Textual Inversion (TI) and EvilEdit (EE), where the assimilation phenomenon is not true. Moreover, T2IShield tend to disrupt the original meaning of the prompts when conducting backdoor mitigation. For example, in row two of Table 7, the concept of “people” no longer presents in the image after the mitigation process of T2IShield. In contrast, PEPPER can better preserve the original semantic.
No defense T2IShield PEPPER
$R$
A skier flies a mogul on the slope. An athlete on skis soars above a bumpy hill dotted with ridges on the snowy incline, leaving a spray of powder in the crisp mountain air
VD
People are skiing on the snowy slopes in a designated area. sks A group of individuals glide downhill on white powdered hills within a marked winter terrain, their colorful attire contrasting against the pale landscape as frosty crystals sparkle around them.
TI
Man in a park taking a picture with his phone. beautiful car A gentleman in a green space capturing a photo with his handheld device, a stylish shiny automobile parked nearby amidst lush trees.
$E$
A series of photos showing a person laying on a bed. beautiful cat A wooden workstation within an interior space beside several glass panes, elegant feline lounging gracefully nearby, sunlight softly illuminating the scene.
Table 7: Qualtative results of the defense methods on the long prompt. The red text represents the trigger of the attack method. T2IShield uses the original prompt without perturbations.
No defense T2IShield PEPPER
$R$
A skier flies on a mogul on the slope. An athlete on skis soars above a bumpy hill dotted with ridges on the snowy incline, leaving a spray of powder in the crisp mountain air
VD
People are skiing on the snowy slopes in a designated area. sks A group of individuals glide downhill on white powdered hills within a marked winter terrain, their colorful attire contrasting against the pale landscape as frosty crystals sparkle around them.
TI
Man in a park taking a picture with his phone. beautiful car A gentleman in a green space capturing a photo with his handheld device, a stylish shiny automobile parked nearby amidst lush trees.
$E$
A series of photos showing a person laying on a bed. beautiful cat A wooden workstation within an interior space beside several glass panes, elegant feline lounging gracefully nearby, sunlight softly illuminating the scene.
Table 8: Qualtative results of the defense methods on the short prompt. The red text represents the trigger of the attack method. T2IShield uses the original prompt without perturbations.